If Sin theta and Cos theta are the roots of the equation ax2+bx+c=0. Then prove that a2-b2+2ac=0

Grade:10

1 Answers

Meet
6 years ago
Let theta be @From the relation of roots sin@+cos@=(-b/a), sin@.cos@=(c/a)Now do squaring both side in first equation and put the values then u will prove that a^2+2ac=b^2.